bewellplus.gsu.edu
Anik4 Anik4
  • 08-11-2015
  • Mathematics
Answered

If the probability of a win is 0.5 and the probability of a draw is 0.3, what is the probability of a loss?

Verified Answer :

Georgie24 Georgie24
  • 08-11-2015

the probability of loss is 0.2 because they must add up to 1.

Answer Link
averageeli averageeli
  • 08-11-2015

0.5+0.3=0.8
1-0.8=0.2
Probability of a loss is 0.2

Answer Link

Related Questions

What effect does screen resolution have on how graphics are displayed
If the sales tax is 85 about how much will an 8995 pair of sneakers cost including sales tax
May you please help me mate
Which of the following decimals is equivalent to 3 and 10 hundreds 030 031 301 or 310
What is 18 as a terminating decimal
Which is usually the best way to present or communicate inferred data in a bar graph in a data table in a simple diagram in a written paragraph
In the term static electric charges what does static refer to
Which of these rocks forms under the MOST extreme conditions of heat and pressure Select one a gneiss b limestone c shale d sandstone
Kacey and her three friends went out for lunch and they wanted to leave a 15 tip The receipt shown below lists the lunch total before tax and tip The tip is on
the soccer team is practicing in small groups there are 36 players in the team how many groups will they have if they split into groups